Two vehicles are travelling towards each other, with a distance of 141 km between them. One is travelling at 70 km/h and the other at 24 km/h. How long will it take for them to meet? Pulling our hair out!!! ;-)
Thank you!!!

Answer by rfer(16322)

approach speed is 70+24=94
t=D/r
t=141/94
t=1.5 hrs